OpenStudy (anonymous):

Prove: if x>=0 and x<=epsilon for all epsilon>0, then x=0.

OpenStudy (anonymous):

@wio would you be able to help with this?

OpenStudy (anonymous):

Hmmm, well what are you supposed to use to do this proof?

OpenStudy (anonymous):

Looks similar to squeeze theorem.

OpenStudy (anonymous):

im not sure what the squeeze theorem is. But I thought maybe you would just use contradiction and assume that either x>0 and x<0, and do both case so then by contradiction x>=0 and x<=0 so that means x=0. but Im really not sure.

OpenStudy (anonymous):

Hmm

OpenStudy (anonymous):

How about something like this... assume x>0. and we must show epsilson>0 such x>epsilon. Let epsilon=x/2. Since x>0, epsilson>0. then epsilson = x/2 = x/2< (x+x)/2 = x. so by contradiction x<=0. or something like that.

OpenStudy (anonymous):

are you still need help?

OpenStudy (anonymous):

yeah, i really dont know what I am doing for if I am even on the right track

OpenStudy (anonymous):

ok,maybe i can help you;)

OpenStudy (ybarrap):

Seems like we need the completeness axiom

OpenStudy (anonymous):

A = {x|x >= 0} B = {x|x <= (epsilon for all epsilon>0)} \[A \lceil \rceil B = {0}\] so, x = 0

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

are A and B like two different cases or are A and B like just different sets.

OpenStudy (anonymous):

I can prove it in this way too: A = {x|x >= 0 , x <= (epsilon for all epsilon>0)} A = {0} so, x = 0

OpenStudy (anonymous):

But if you are given that x>=0 and e>0. then all you are saying is that A=0 so x must equal 0. that is all you have to say??

OpenStudy (anonymous):

yes ;)

OpenStudy (anonymous):

easy enough. thank you.

OpenStudy (anonymous):

your welcome :)

OpenStudy (zzr0ck3r):

density of rational/irrationals..

OpenStudy (zzr0ck3r):

assume \[x\ge0,x\le\epsilon\text{ for all }\epsilon>0\\\text{assume by contradiction }x\ne0\\\text{by density we have that there exists }\epsilon \le x\\a \space contradiction...\]